The Thompson Corporation, a manufacturer of steel products, began operations on October 1, 2019. The accounting department of Thompson has started the fixed-asset and depreciation schedule presented below. You have been asked to assist in completing this schedule. In addition to ascertaining that the data already on the schedule are correct, you have obtained the following information from the company's records and personnel: (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

  1. Depreciation is computed from the first of the month of acquisition to the first of the month of disposition.
  2. Land A and Building A were acquired from a predecessor corporation. Thompson paid $742,500 for the land and building together. At the time of acquisition, the land had a fair value of $99,600 and the building had a fair value of $730,400.
  3. Land B was acquired on October 2, 2019, in exchange for 2,300 newly issued shares of Thompson’s common stock. At the date of acquisition, the stock had a par value of $5 per share and a fair value of $18 per share. During October 2019, Thompson paid $9,700 to demolish an existing building on this land so it could construct a new building.
  4. Construction of Building B on the newly acquired land began on October 1, 2020. By September 30, 2021, Thompson had paid $140,000 of the estimated total construction costs of $230,000. Estimated completion and occupancy are July 2022.
  5. Certain equipment was donated to the corporation by the city. An independent appraisal of the equipment when donated placed the fair value at $13,200 and the residual value at $1,300.
  6. Equipment A’s total cost of $115,000 includes installation charges of $480 and normal repairs and maintenance of $10,300. Residual value is estimated at $5,700. Equipment A was sold on February 1, 2021.
  7. On October 1, 2020, Equipment B was acquired with a down payment of $3,300 and the remaining payments to be made in 10 annual installments of $3,300 each beginning October 1, 2021. The prevailing interest rate was 7%.


Required:

Supply the correct amount for each answer box on the schedule. (Round your intermediate calculations and final answers to the nearest whole dollar.)THOMPSON CORPORATION Fixed Asset and Depreciation Schedule For Fiscal Years Ended September 30, 2020, and September 30, 2021
Answer #1

Solution Given Data Thompson paid $742,500 to the land and building together. Alt the time of acquired sition, the land had aBuilding A cost $749500 x 790400 99600 + 730400 = 742500 * 730400 Sam X100) 830,000 = 742500 x 7804001 8300 = 742500 * 887 =Building B De pre ü ation for the year ended 9/30/2021 will not be change before it is as under construction, Donated equipme• qa430 , Sy 49480 x+63-0.05454 = $5426 Equipment Bi Depreciation to? yecon ended 9/30/2021 1- (1+77) 17 3300 X 7 3300 X 7.51Assets Acquistion date . De prolon 407 year ended a 30 cost | Residual De par ciddion method estimated like invel 2020 2021 l